OpenStudy (anonymous):

4 - 7x + 10 = 15 - 6x - 8

OpenStudy (uri):

-7x+6x=15-8-10-4 - x=-7 x=7 ans @bonillahenry18

OpenStudy (anonymous):

gracias

OpenStudy (uri):

:3